OpenStudy (anonymous):

A jar contains 24 black ad 6 purple marbles. A marble is drawn at random. P(black) so would that be 50/24

jimthompson5910 (jim_thompson5910):

there are 24 black marbles and 30 marbles total

jimthompson5910 (jim_thompson5910):

so P(Black) = 24/30 P(Black) = 5/6

OpenStudy (anonymous):

24/30 simplifyes 12/15

OpenStudy (anonymous):

simplified*

OpenStudy (anonymous):

then 4/5

OpenStudy (anonymous):

Lol i calculated the wrong number thanks

jimthompson5910 (jim_thompson5910):

oops made a typo P(Black) = 24/30 P(Black) = 4/5 yeah Guy caught it before I did lol

OpenStudy (anonymous):

lol